Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Thu Jun 4 23:22:39 METDST 1998) Success
Experiment successfully carried out. No obvious problems reported.

Westerbork (Tony Foley at Fri Jun 5 09:36:05 METDST 1998) Minor failures
One BIG problem, we used a thick tape to start with so all the stuff before we switched to thin tape (at 1847UT) will probably be useless. Data from 1930 to 2130 should be OK, despite the warning of bad parity errors (the standard check procedures still assume low density recording).

Torun (K. Borkowski at Fri Jun 5 10:20:48 METDST 1998) Success
We have used a THIN tape, despite a clear remark in the c98c2lst file to the contrary. Hopefully it is what should really be.

Onsala (John Conway at Fri Jun 5 14:14:28 METDST 1998) Minor failures
Finshed experiment early (around 21:15UT), due to failure of head positioning, subsequently traced to the MAT card which was replaced. No other known problems. Tape shipped about midday Froday by DHL to Bonn.

Jodrell (Alastair Gunn at Sun Jun 7 18:16:26 METDST 1998) Success
No known problems.

Cambridge (Alastair Gunn at Sun Jun 7 18:16:58 METDST 1998) Success
No known problems.

Noto (C.Trigilio at Mon Jun 8 17:37:43 METDST 1998) Success

Urumqi (Liu Xiang at Wed Jun 10 11:12:24 METDST 1998) Success
The Clock Offset is 1.7us. It is rain and cloud. The tape is sent in 10, June.

Calibration (Jean\-Francois Desmurs at Wed Jun 10 15:55:30 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the ? directory under the name c98l2.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 14:52:15 METDST 1998) Minor failures
off source 1730 to 1743